OpenStudy (anonymous):

Sand is being poured on the ground at the rate of 25m^3/min forming a conical pile whose altitude is the same as the radius of the base. How fast is the altitude increasing when the diameter of the base is 12m?

OpenStudy (nincompoop):

set up your equation

OpenStudy (anonymous):

V= 1/3pi(r)^2h

OpenStudy (nincompoop):

and what is the question?

OpenStudy (anonymous):

I'm not sure how to find the rate of the altitude?

OpenStudy (nincompoop):

25m^3/min forming a conical pile whose altitude is the same as the radius of the base ow fast is the altitude increasing when the diameter of the base is 12m?

OpenStudy (nincompoop):

\(V'(t) = 25 \) and you need \(r(t) \) = \(h(t)\)

OpenStudy (nincompoop):

\(r(t) = \frac{d}{2} = 6\)

OpenStudy (anonymous):

d=12

OpenStudy (nincompoop):

ye, but your equation for volume uses radius and not diameter

OpenStudy (anonymous):

radius and the height are both 6m

OpenStudy (nincompoop):

ye

OpenStudy (anonymous):

do I sub the 6 into the equation for both radius and height?

OpenStudy (nincompoop):

so you're solving for \(\huge \frac{dh}{dt}\)

OpenStudy (nincompoop):

but you need \(\large \frac{dV}{dt} =\)

OpenStudy (anonymous):

25

OpenStudy (nincompoop):

equation

OpenStudy (nincompoop):

then put the value

OpenStudy (nincompoop):

how about this, try to draw what is going on

OpenStudy (anonymous):

v=1/3pi(6^2(h)) dv/dt=12pi[dh/dt]

OpenStudy (nincompoop):

I think it will help you if you can visually interpret it

OpenStudy (nincompoop):

the reason why I am asking is because you're given a diameter, which tells you the area of the base

OpenStudy (anonymous):

Im not sure how to draw it, but I could try...